A chemistry student has a 15% hydrochloric acid solution and a 40% hydrochloric acid solution. How many liters of each should she mix to get 10 liters of a hydrochloric acid solution with a 30% acid concentration?
Answer by Maths68(1474)

Solution A
Amount = x liters
Concentration =15% = 0.15
=============================================
Solution B
Amount =(10- x) liters
Concentration =40% = 0.40
=============================================
Resultant Solution
Amount = 10 liters
Concentration = 30%=0.30
=============================================
[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B] = Amount of Resultant * Concentration of resultant
==============================================
Put the value in above equation
(x*0.15)+((10-x)*0.40=10*0.30
0.15x+4-0.4x=3
-0.25x=3-4
-0.25x=-1
Divide by -0.25 both sides
-0.25x/-0.25=-1/-0.25
x=4
=============================================
Amount of solution A =4 liters
Amount of Sultion B = 10-x=10-4=6 liters
==============================================
A chemistry student has a 15% hydrochloric acid solution and a 40% hydrochloric acid solution.He has to mix 4 liters of 15% solution and 6 liters of 40% to get 10 liters of a hydrochloric acid solution with a 30% acid concentration?